OpenStudy (anonymous):

Please Help!!!!! Fan and medal!!!!! Which of these ordered pairs is a solution to the inequality y – 2x ≤ –3? A. (2, 4) B. (–2, 3) C. (1, –1) D. (3, 4)

OpenStudy (anonymous):

@texaschic101

OpenStudy (sweetburger):

plug them in and then tell me which one you think it is

OpenStudy (anonymous):

how

OpenStudy (sweetburger):

you have a y and a x value and you have coordinates that represent (x,y) plug those in and see which ones work

OpenStudy (anonymous):

i dont get it

OpenStudy (texaschic101):

each of your answer choices are in (x,y) form y - 2x <= -3 take your first answer choice...(2,4)..x = 2 and y = 4 now sub 2 in for x and 4 in for y into your equation to see if its true 4 - 2(2) <= -3 4 - 4 <= -3 0 <= -3 now is 0 less then or equal to -3. No it is not. Therefore, it is not the correct answer understand ? Now just sub your other answer choices into the equation and whichever one is true, is the correct answer

OpenStudy (texaschic101):

can there be more then 1 answer ?

OpenStudy (anonymous):

no

OpenStudy (texaschic101):

do you have the inequality sign written correctly ?

OpenStudy (anonymous):

yes

OpenStudy (texaschic101):

(1,-1) y - 2x <= -3 -1 - 2(1) <= -3 -1 - 2 <= -3 -3 <= -3 (correct) -- this is an answer (3,4) y - 2x <= -3 4 - 2(3) <= -3 4 - 6 <= -3 -2 <= -3 (correct) -- this is an answer first one is not...second one is not....you have 2 answers

OpenStudy (texaschic101):

wait..I messed up...I was looking at the problem wrong.....answer is (1,-1) not (3,4)...sorry

OpenStudy (texaschic101):

because with (3,4).....-2 is not less then -3...it is greater

OpenStudy (anonymous):

can you help me with two more please

OpenStudy (texaschic101):

sure

OpenStudy (texaschic101):

x <= 2.....when you have x = a number, then the line is vertical. There is an equal sign in the problem, meaning the line is solid. Less then means shading to the left. your correct graph is : A

OpenStudy (anonymous):

there mixed up options so option D

OpenStudy (texaschic101):

yes..option D

OpenStudy (anonymous):

I have one more

OpenStudy (texaschic101):

lets have it :)

OpenStudy (texaschic101):

3y - 5x >= -6 3y = 5x - 6 y = 5/3 - 2 slope is 5/3 y int = (0,-2) x int : 3(0) - 5x = -6 -5x = -6 x = 6/5 or for graphing purposes, x = 1 1/5 solid line because of equal sign greater then means shading above the line graph is : option A

OpenStudy (anonymous):

100% thank you thank you so much